# Key Words.: Gravitational lensing: strong – Methods: data analysis – Methods: observational – Galaxies: clusters: individual: Perseus1 Introduction

### v1

@@
We investigate the ability of the Euclid telescope to detect galaxy-scale gravitational lenses. To do so, we perform a systematic visual inspection of the 0.7 deg2 Euclid ERO data towards the Perseus cluster using both the high-resolution VIS band, and the lower resolution NISP bands. We inspect every extended source brighter than magnitude 23 in with 41 expert human classifiers. This amounts to 12 086 stamps of . We find 3 grade A and 13 grade B candidates. We assess the validity of these 16 candidates by modelling them and checking that they are consistent with a single source lensed by a plausible mass distribution. Five of the candidates pass this check, five others are rejected by the modelling and six are inconclusive. Extrapolating from the five successfully modelled candidates, we infer that the full of the Euclid Wide Survey should contain galaxy-galaxy lenses that are both discoverable through visual inspection and have valid lens models. This is consistent with theoretical forecasts of discoverable galaxy-galaxy lenses in Euclid. Our five modelled lenses have Einstein radii in the range , but their Einstein radius distribution is on the higher side when compared to theoretical forecasts. This suggests that our methodology is likely missing small Einstein radius systems. Whilst it is implausible to visually inspect the full Euclid data set, our results corroborate the promise that Euclid will ultimately deliver a sample of around galaxy-scale lenses.
@@
Strong gravitational lensing by massive galaxies offers a plethora of applications both in cosmology and astrophysics. Some notable examples include measuring the total mass of lens galaxies within the Einstein radius, as well as disentangling the contributions from visible and dark matter components (e.g., Auger et al. 2009). When coupled with deep spectroscopic observations, it allows for the placement of constraints on the stellar initial mass function of lens galaxies (e.g., Ferreras et al. 2010; Dutton & Treu 2014; Sonnenfeld et al. 2019). Thanks to the lensing magnification, it serves as a natural telescope to study lensed sources otherwise too faint or angularly too small to be detected (e.g., Hezaveh et al. 2013), and even to map their velocity field with unprecedented spatial resolution (e.g., Paraficz et al. 2018). By studying small-scale distortions of lensed images or arcs, it allows us to infer the presence of low-mass dark halos either in the lens or along the line of sight up to the source redshift, and in turn to study the properties of dark matter (e.g., Vegetti et al. 2010; O’Riordan et al. 2023; Gilman et al. 2024). Moreover, when the lensed source is time-variable, such as a quasar or a supernova, strong lensing offers an independent way of measuring the expansion rate of the Universe, (Refsdal 1964; Wong et al. 2020; Shajib et al. 2023; Grillo et al. 2024; Pascale et al. 2024).
@@
Lastly, the measurement of weak lensing shear from strong lensing images has been proposed as a potential cosmological probe (Birrer et al. 2017, 2018). A minimal model for the shear that is non-degenerate with lens model parameters was derived by Fleury et al. (2021), and this quantity was shown to be measurable in mock imaging data by Hogg et al. (2023). For high-precision cosmological constraints, however, a Euclid-size data set of strong lenses will be required.
@@
Galaxy-scale strong lensing events are rare, with about one object out of thousands showing lensing features (Oguri & Marshall 2010; Collett 2015). In addition, these systems are small in angular terms, spanning between tenths of an arcsecond to a few arcseconds on the sky, and the lensed images of the sources (arcs, rings, and multiple point sources) are often hidden in the glare of the foreground lensing galaxy. Because lenses are rare, compact, and low-contrast objects, they are best discovered in deep, sharp, wide-field surveys. Euclid is exactly that (Euclid Collaboration: Mellier et al. 2024), and is the focus of the present work. In fact, lensed QSO candidates have already been proposed from Euclid observations (Cuillandre et al. 2024b).
@@
However, even with high-quality Euclid data, finding lenses remains challenging, not only because of the intrinsic rarity of strong lensing, but also because other non-lensing objects mimic the morphology of lenses (e.g., ring galaxies, galaxy mergers, spiral galaxies, or even random alignment). So far, the best methods to discover lenses involve various flavours of convolutional neural networks (CNNs; among many others see Petrillo et al. 2019; Jacobs et al. 2019; Stein et al. 2022; Savary et al. 2022; Rojas et al. 2022) and more recently, transformer networks (Thuruthipilly et al. 2022; Grespan et al. 2024). These networks require training sets that match the instrumental and astrophysical properties of the data very closely. Since we do not possess a number of real strong lens observations sufficient for a training set, these have to be simulated, and the effectiveness of simulations in emulating nature is limited.
@@
Machine learning methods are effective in finding lenses, but they are limited in terms of purity and false-positive rates. When false-positive rates are below 1% (e.g., Canameras et al. 2023), they are useful to pre-select candidates among parent samples of tens or even hundreds of millions of galaxies. However, as the true – and poorly known – prevalence of lenses on the plane of the sky is very low, even a 1% false-positive rate can result in samples dominated by contaminants and unconvincing candidates. This is known as the base rate fallacy. As a result, human visual inspection is almost always necessary to fine-grain the network selection, although humans themselves sometimes have difficulty deciding on the validity of a lens candidate. This is especially true for small Einstein radius lenses.
@@
Human visual inspection is useful on its own as a lens finding method, but is limited by the volume of data that can be inspected by a team of experts in a reasonable amount of time. However, it has the advantage that it can pick up unusual lensing configurations (e.g., Keeton et al. 2000; Orban de Xivry & Marshall 2009; Collett & Bacon 2016) that, by definition, do not appear in large numbers in simulated training sets. It is also the only way to evaluate the prevalence of lenses on the plane of the sky and therefore the expected surface density of lensing systems for a given instrument, depth, spatial resolution, and wavelength. This has been attempted a few times in the past, for example in Hubble Space Telescope (HST) imaging (e.g., Faure et al. 2008; Pawase et al. 2014; Garvin et al. 2022). Rojas et al. (2023) also evaluate the performance of visual inspection with simulated data.
@@
In the Early Release Observations (ERO) Lens Search Experiment (ELSE), we carry out a blind visual search of galaxy-scale strong lensing systems, using some of the first data from the Euclid ESA satellite (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024). We focus on the Euclid ERO imaging of the Perseus galaxy cluster (Cuillandre et al. 2024a) and the only selection criterion for the extended sources to be inspected is a magnitude cut. Our search is therefore one of the broadest visual searches carried out so far in terms of pre-selection. The goals are: (1) to evaluate the performance of the Euclid telescope at finding galaxy-scale lenses; (2) to study the prevalence of strong lenses found by humans in Euclid; (3) to test the efficiency of human experts at finding lenses; and, (4) to optimise future visual inspections that will be performed on lens candidates found by the automatic pipeline in Euclid.
@@
As the humans involved in this exercise are all lensing experts but limited in number (41 exactly), our work contrasts with more intensive citizen science searches involving much larger numbers of humans but spanning a much scarcer range in terms of expertise available (Marshall et al. 2015; More et al. 2016).
@@
This article is organised as follows. We introduce the Euclid ERO observations of the Perseus cluster in Sect. 2. Then, we explain the methodology of the visual inspections in Sect. 3, along with the visualisation tools used for them. Afterwards, we report the results from the visual inspection in Sect. 4. Next, we present the sample of lens candidates and some initial modelling in Sect. 5. Then, in Sect. 6 we compare the sample against the literature and estimate the number of lenses that will be discovered in the Euclid Wide Survey (EWS). Finally, we conclude in Sect. 7.
@@ 2 Data
The Euclid ERO programme (Euclid Early Release Observations 2024) targeted the Perseus galaxy cluster, obtaining very deep data of the central region of the cluster in , in the broad optical filter from the visible imager (VIS) instrument (Euclid Collaboration: Cropper et al. 2024), and the three broad filters , , and in the near-infrared from the Near-Infrared Spectrometer and Photometer (NISP) instrument (Euclid Collaboration: Jahnke et al. 2024).
These data were collected during the Euclid performance verification phase in September 2023 (Cuillandre et al. 2024a). All Euclid science observations adhere to a reference observing sequences (ROS, Euclid Collaboration: Scaramella et al. 2022), which consists of four dithered exposures lasting seconds each in the filter, and four dithered exposures of seconds each in the , , and filters. Four ROSs were obtained for this field, with a total integration time of seconds in the filter and seconds in the , , and filters, achieving a depth magnitudes deeper than that of the EWS, which relies on a single ROS (Euclid Collaboration: Scaramella et al. 2022). Therefore, these exceptional data reach a point-source depth of () at , with a () full width at half maximum (FWHM), and a surface brightness limit of . We refer the reader to Cuillandre et al. (2024a) for more details on the data reduction.
The Perseus cluster data set represents the ideal data to perform a first search of gravitational lenses, and prepare for the future systematic investigations with the EWS data.
The astrometrically and photometrically calibrated imaging stacks across all four Euclid bands are accompanied by catalogues for compact sources produced using the tool SourceExtractor (Bertin & Arnouts 1996) on the “flattened” stack, i.e., with the low spatial frequencies removed (sky background and any Galactic cirrus or nebulae). Multi-wavelength catalogues including VIS, NISP, and ground-based CFHT-MegaCam photometry have been provided in Cuillandre et al. (2024b). After star-galaxy separation and limiting the analysis to sources with , we end up with a parent sample of objects, which will be used in the visual classification procedure. As shown in Fig. 1, the adopted magnitude cut is well motivated by the fact that sources at fainter magnitudes are compact and featureless.
For the experiment, cutouts in the four bands of (i.e. pixels and pixels in the VIS and NISP bands, respectively) have been created using the “flattened” stacks.
@@ 3.1 The visualisation tools
We use the visualisation tools prepared by Acevedo Barroso et al. (in prep.) to carry out the visual inspection.111The version used for this work is available at https://github.com/ClarkGuilty/Qt-stamp-visualizer/tree/ERO_edition The tools are based on the visualisation tools used in Savary et al. (2022) and Rojas et al. (2022), but reimplemented in the Qt 6 framework with extended functionality targeting the requirements of the Euclid ERO data. The two applications correspond to a mosaic viewer and a 1-by-1 sequential viewer, both presented in Fig. 2. The mosaic viewer displays rectangular mosaics of objects for the user to inspect. Then, the user is tasked with clicking on objects that show any signs of lensing. Additionally, the user is allowed to mark objects as “interesting, but not a lens”. This is exemplified in the left panel of Fig. 2. By contrast, the 1-by-1 sequential tool displays one object at the time and the user is tasked with classifying it into one of the following non-overlapping categories:
A: sure lens, it shows clear lensing features and no additional information is needed;
B: probable lens, it shows lensing features but additional information is required to verify that it is a definite lens;
C: possible lens, it shows lensing features, but they can be explained without resorting to gravitational lensing;
X: definitively not a lens;
Interesting: definitively not a lens, but interesting in some other way.
The grades A, B, and C are deemed as positive grades, whereas X and Interesting are deemed as negative. The 1-by-1 sequential tool is shown in the right panel of Fig. 2. Both tools allow the users to change the colour map used for the monochromatic images, as well as the function used to scale the pixel intensities. For every object to inspect, the tools generate three high-contrast images:
@@
an RGB composite image using the , , and bands, at the NISP resolution of ;
@@
Before creating the composite images, we reproject the band data from the sky-coordinate system and resolution of the VIS instrument, to the corresponding ones in the NISP instrument. This aligns the images and corrects for the different pixel scales between instruments. The experts were required to use the monochromatic high-resolution band, and the composite. The composite is not shown by default in the tools, but it is also available. Additionally, the users also have access to the individual NISP bands when using the 1-by-1 sequential tool. This is shown in Fig. 2.
@@ 3.2 The visual inspections
To make better use of the large number of strong lensing experts available for the visual inspection, we split the 41 experts into two non-overlapping teams. This allows us to try different visual inspection schemes while still retaining enough experts per team for the classifications not to be dominated by noise. Team 1 comprises 22 experts, while Team 2 comprises 19. Both teams inspect the cutouts described in Sect. 2, along with six extra cutouts of simulated lenses produced following the prescription of Euclid Consortium, Metcalf et al. (in prep). The introduction of simulated mocks tests the ability of the experts to identify lenses, and serves as a sanity check of their performance. The experts are not informed about the mocks in order to avoid any biases. The six simulated lenses can be seen in Fig. 3. Both teams are given two weeks per stage. There is no communication either between or within the teams and their results are not joined until after the visual inspections are completed.
Both teams follow a two-stage approach. First, they focus on cleaning the parent sample from obvious non-lens contaminants. Then, they inspect in detail the remaining sources and settle on a final classification following the grades introduced in Sect. 3.1.
Team 1 first inspects the entire parent sample using the mosaic tool. This corresponds to cutouts after adding the mock lenses. The experts are allowed to change the number of sources per page shown in the mosaic, with most experts observing between 20 and 42 sources per mosaic page. Afterwards, Team 1 reinspects all the objects selected in the first stage using the 1-by-1 sequential tool and assigning detailed classifications.
Simultaneously, Team 2 is further divided into three groups: two groups comprising six experts each and one group comprising seven. Each group inspects one third of the sources using the 1-by-1 sequential tool, while focusing on rejecting obvious contaminants. During the first stage, no detailed classification is required. Subsequently, all the selected sources are reinspected by Team 2 as a whole, using again the 1-by-1 sequential tool but assigning detailed classifications.
Afterwards, we aggregate the final classifications of each team independently. The output is a single final classification per team for every source. Thus, if more than half of the experts voted negatively, then the source is taken as a non-lens; otherwise, we take the mode between A, B and C. In the case of a tie between A and B or A and C, A takes precedence. This prioritisation reflects the expectation that experts will only vote for A if they are confident about the presence of lensing features. If there is a tie between B and C, we favour C to minimise noise in the final B-sample. In the event of a tie among all three positive classes, B is preferred.

### v2

@@
We investigated the ability of the Euclid telescope to detect galaxy-scale gravitational lenses. To do so, we performed a systematic visual inspection of the 0.7 deg2 Euclid Early Release Observations data towards the Perseus cluster using both the high-resolution band and the lower-resolution , , and bands. Each extended source brighter than magnitude 23 in was inspected by 41 expert human classifiers. This amounts to 12 086 stamps of . We found 3 grade A and 13 grade B candidates. We assessed the validity of these 16 candidates by modelling them and checking that they are consistent with a single source lensed by a plausible mass distribution. Five of the candidates pass this check, five others are rejected by the modelling, and six are inconclusive. Extrapolating from the five successfully modelled candidates, we infer that the full of the Euclid Wide Survey should contain galaxy-galaxy lenses that are both discoverable through visual inspection and have valid lens models. This is consistent with theoretical forecasts of discoverable galaxy-galaxy lenses in Euclid. Our five modelled lenses have Einstein radii in the range , but their Einstein radius distribution is on the higher side when compared to theoretical forecasts. This suggests that our methodology is likely missing small-Einstein-radius systems. Whilst it is implausible to visually inspect the full Euclid dataset, our results corroborate the promise that Euclid will ultimately deliver a sample of around galaxy-scale lenses.
@@
Strong gravitational lensing by massive galaxies offers a plethora of applications in both cosmology and astrophysics. Some notable examples include measuring the total mass of lens galaxies within the Einstein radius and disentangling the contributions from visible and dark matter components (e.g. Auger et al., 2009). When coupled with deep spectroscopic observations, it enables the placement of constraints on the stellar initial mass function of lens galaxies (e.g. Ferreras et al., 2010; Dutton & Treu, 2014; Sonnenfeld et al., 2019). Thanks to the lensing magnification, strong gravitational lensing serves as a natural telescope to study lensed sources otherwise too faint or angularly too small to be detected (e.g. Hezaveh et al., 2013), and even to map their velocity field with unprecedented spatial resolution (e.g. Paraficz et al., 2018). Studying small-scale distortions of lensed images or arcs allows us to infer the presence of low-mass dark halos either in the lens or along the line of sight up to the source redshift, and in turn to study the properties of dark matter (e.g. Vegetti et al., 2010; O’Riordan et al., 2023; Gilman et al., 2024). Moreover, when the lensed source is time-variable, such as a quasar or a supernova, strong lensing offers an independent way of measuring the expansion rate of the Universe, (Refsdal, 1964; Wong et al., 2020; Shajib et al., 2023; Grillo et al., 2024; Pascale et al., 2025).
@@
Lastly, the measurement of weak lensing shear from strong lensing images has been proposed as a potential cosmological probe (Birrer et al., 2017, 2018). A minimal model for the shear that is non-degenerate with lens model parameters was derived by Fleury et al. (2021), and this quantity was shown to be measurable in mock imaging data by Hogg et al. (2023). For high-precision cosmological constraints, however, a Euclid-sized dataset of strong lenses will be required.
@@
Galaxy-scale strong lensing events are rare, with about one object out of thousands showing lensing features (Oguri & Marshall, 2010; Collett, 2015). In addition, these systems are small in angular terms, spanning between tenths of an arcsecond to a few arcseconds on the sky, and the lensed images of the sources (arcs, rings, and multiple point sources) are often hidden in the glare of the foreground lensing galaxy. Because lenses are rare, compact, and low-contrast objects, they are best discovered in deep, sharp, wide-field surveys. Euclid is exactly that (Euclid Collaboration: Mellier et al., 2025), and is the focus of the present work. In fact, lensed quasi-stellar object candidates have already been proposed from Euclid observations (Cuillandre et al., 2025b).
@@
However, even with high-quality Euclid data, finding lenses remains challenging, not only because of the intrinsic rarity of strong lensing, but also because other non-lensing objects mimic the morphology of lenses (e.g. ring galaxies, galaxy mergers, spiral galaxies, or even random alignments). So far, the best methods for discovering lenses involve various flavours of convolutional neural networks (see Petrillo et al. 2019; Jacobs et al. 2019; Stein et al. 2022; Savary et al. 2022; Rojas et al. 2022, to cite just a few) and, more recently, transformer networks (Thuruthipilly et al., 2022; Grespan et al., 2024; González et al., 2025). These networks require training sets that match the instrumental and astrophysical properties of the data very closely. Since we do not possess enough strong lens observations to constitute a training set, they have to be simulated, and the effectiveness of simulations in emulating nature is limited.
@@
Machine learning methods are effective at finding lenses, but they are limited in terms of purity and false-positive rates. When false-positive rates are below 1%, such methods are useful for pre-selecting candidates from parent samples of tens or even hundreds of millions of galaxies (Cañameras et al., 2024). However, as the true – and poorly known – prevalence of lenses on the plane of the sky is very low, even a 1% false-positive rate can result in samples dominated by contaminants and unconvincing candidates. This is known as the base rate fallacy. As a result, human visual inspection is almost always necessary to fine-tune the network selection, although humans themselves sometimes have difficulty deciding on the validity of a lens candidate. This is especially true for small-Einstein-radius lenses.
@@
Human visual inspection is useful on its own as a lens finding method but is limited by the volume of data that can be inspected by a team of experts in a reasonable amount of time. However, it has the advantage that it can pick up unusual lensing configurations (e.g. Keeton et al., 2000; Orban de Xivry & Marshall, 2009; Collett & Bacon, 2016) that, by definition, do not appear in large numbers in simulated training sets. It is also the only way to evaluate the prevalence of lenses on the plane of the sky and therefore the expected surface density of lensing systems for a given instrument, depth, spatial resolution, and wavelength. This has been attempted a few times in the past, for example in Hubble Space Telescope imaging (e.g. Faure et al., 2008; Pawase et al., 2014; Garvin et al., 2022). Rojas et al. (2023) also evaluated the performance of visual inspection with simulated data.
@@
In the Early Release Observations (ERO) Lens Search Experiment (ELSE), we carried out a blind visual search of galaxy-scale strong lensing systems using some of the first data from the European Space Agency (ESA) satellite Euclid (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025). We focused on the Euclid ERO imaging of the Perseus galaxy cluster (Cuillandre et al. 2025a), and the only selection criterion for the extended sources to be inspected was a magnitude cut. Our search is therefore one of the broadest visual searches carried out so far in terms of pre-selection. The goals were: (1) to evaluate the performance of the Euclid telescope at finding galaxy-scale lenses; (2) to study the prevalence of strong lenses found by humans in Euclid; (3) to test the efficiency of human experts at finding lenses; and (4) to optimise future visual inspections that will be performed on lens candidates found by the automatic pipeline in Euclid.
@@
The humans involved in this exercise are all lensing experts but were limited in number (41 exactly). Our work therefore contrasts with more intensive citizen science searches involving much larger numbers of humans but spanning a much smaller range in terms of expertise (Marshall et al., 2015; More et al., 2016).
Finally, assuming subtraction of the lens light, Euclid is predicted to find lenses in its six-year main survey (Collett, 2015). We tested that prediction in this work, with the caveat that we did not perform any subtraction of the lens light. Thus, our results represent a lower bound in this regard.
@@
We introduce the Euclid ERO observations of the Perseus cluster in Sect. 2. We then explain the methodology of the visual inspections in Sect. 3, along with the visualisation tools used. Afterwards, we report the results from the visual inspection in Sect. 4 and present the sample of lens candidates and some initial modelling in Sect. 5. In Sect. 6 we then compare the sample against the literature and estimate the number of lenses that will be discovered in the Euclid Wide Survey (EWS). We conclude in Sect. 7.
@@ 2 Data
The Euclid ERO programme (Euclid Early Release Observations, 2024) targeted the Perseus galaxy cluster, obtaining very deep data of the central region of the cluster in , in the broad optical filter from the visible imager (VIS) instrument (Euclid Collaboration: Cropper et al., 2025), and the three broad filters , , and in the near-infrared from the Near-Infrared Spectrometer and Photometer (NISP) instrument (Euclid Collaboration: Jahnke et al., 2025).
These data were collected during the Euclid performance verification phase in September 2023 (Cuillandre et al., 2025a). All Euclid science observations adhere to a reference observing sequences (ROSs; Euclid Collaboration: Scaramella et al., 2022), which consists of four dithered exposures lasting seconds each in the filter, and four dithered exposures of seconds each in the , , and filters. Four ROSs were obtained for this field, with a total integration time of seconds in the filter and seconds in the , , and filters, achieving a depth magnitudes deeper than that of the EWS, which relies on a single ROS (Euclid Collaboration: Scaramella et al., 2022). Therefore, these exceptional data reach a point-source depth of () at , with a () full width at half maximum, and a surface brightness limit of . We refer the reader to Cuillandre et al. (2025a) for more details on the data reduction.
The astrometrically and photometrically calibrated imaging stacks across all four Euclid bands are accompanied by catalogues for compact sources produced using the tool SourceExtractor (Bertin & Arnouts, 1996) on the ‘flattened’ stack, that is, with the low spatial frequencies removed (sky background and any Galactic cirrus or nebulae). Multi-wavelength catalogues including VIS, NISP, and ground-based Canada-France-Hawaii Telescope MegaCam photometry have been provided in Cuillandre et al. (2025b). After star–galaxy separation and limiting the analysis to sources with , we ended up with a parent sample of objects, which was later used in the visual classification procedure. As shown in Fig. 1, the adopted magnitude cut is justified by the fact that sources at fainter magnitudes are compact and featureless. For the experiment, cutouts in the four Euclid bands of (i.e. pixels and pixels in the VIS and NISP bands, respectively) were created using the flattened stacks.
@@ 3.1 The visualisation tools
We used the visualisation tools prepared by Acevedo Barroso et al. (2025) to carry out the visual inspection.111The version used for this work is available at https://github.com/ClarkGuilty/Qt-stamp-visualizer/tree/ERO_edition The tools are based on the visualisation tools used in Savary et al. (2022) and Rojas et al. (2022), but re-implemented in the Qt 6 framework with extended functionality targeting the requirements of the Euclid ERO data. The two applications correspond to a mosaic viewer and a one-by-one sequential viewer, both presented in Fig. 2. The mosaic viewer displays rectangular mosaics of objects for the user to inspect. Then, the user is tasked with clicking on objects that show any signs of lensing. Additionally, the user is allowed to mark objects as ‘interesting, but not a lens’. This is exemplified in the left panel of Fig. 2. By contrast, the one-by-one sequential tool displays one object at the time and the user is tasked with classifying it into one of the following non-overlapping categories:
‘A’ indicates a sure lens: it shows clear lensing features and no additional information is needed.
‘B’ indicates a probable lens: it shows lensing features but additional information is required to verify it as a definite lens.
‘C’ indicates a possible lens: it shows lensing features, but they can be explained without resorting to gravitational lensing.
‘X’ indicates it is definitively not a lens.
‘Interesting’ indicates it is definitively not a lens but is interesting in some other way.
The grades A, B, and C are deemed as positive grades, whereas X and Interesting are deemed as negative. The one-by-one sequential tool is shown in the right panel of Fig. 2. Both tools allow the users to change the colour map used for the monochromatic images, as well as the function used to scale the pixel intensities. For every object to inspect, the tools generate three high-contrast images:
@@
a red-green-blue (RGB) composite image using the , , and bands, at the NISP resolution of ;
@@
Before creating the composite images, we re-projected the band data from the sky-coordinate system and resolution of the VIS instrument to the corresponding ones in the NISP instrument. This aligns the images and corrects for the different pixel scales between instruments. The experts were required to use the monochromatic high-resolution band, and the composite. The composite is not shown by default in the tools, but it is also available. Additionally, the users also have access to the individual NISP bands when using the one-by-one sequential tool. This is shown in Fig. 2.
@@ 3.2 The visual inspections
To make better use of the large number of strong lensing experts available for the visual inspection, we split the 41 experts into two non-overlapping teams. This allowed us to try different visual inspection schemes while still retaining enough experts per team for the classifications not to be dominated by noise. Team 1 comprises 22 experts, and Team 2 comprises 19. Both teams inspect the cutouts described in Sect. 2, along with six extra cutouts of simulated lenses produced following the prescription of Euclid Consortium, Metcalf et al. (in prep). The introduction of simulated mocks tests the ability of the experts to identify lenses, and serves as a sanity check of their performance. The experts are not informed about the mocks in order to avoid any biases. The six simulated lenses can be seen in Fig. 3. Both teams are given two weeks per stage. There is no communication either between or within the teams, and their results are not combined until after the visual inspections are completed.
Both teams followed a two-stage approach. First, they focused on cleaning the parent sample of obvious non-lens contaminants. Then, they inspected the remaining sources in detail and settled on a final classification following the grades introduced in Sect. 3.1.
Team 1 first inspects the entire parent sample using the mosaic tool. This corresponds to cutouts after adding the mock lenses. The experts were allowed to change the number of sources per page shown in the mosaic, with most experts observing between 20 and 42 sources per mosaic page. Afterwards, Team 1 reinspected all the objects selected in the first stage using the one-by-one sequential tool and assigning detailed classifications.
Simultaneously, Team 2 is further divided into three groups: two groups comprising six experts each and one group comprising seven. Each group inspects one third of the sources using the one-by-one sequential tool, while focusing on rejecting obvious contaminants. During the first stage, no detailed classification is required. Subsequently, all the selected sources are reinspected by Team 2 as a whole, using again the one-by-one sequential tool but assigning detailed classifications.
Afterwards, we aggregated the final classifications of each team independently. The output is a single final classification per team for every source. Thus, if more than half of the experts voted negatively, then the source was classified as a non-lens; otherwise, we classified it as the majority vote between A, B, and C. In the case of a tie between A and B, or A and C, A took precedence. This prioritisation reflects the expectation that experts will only vote for A if they are confident about the presence of lensing features. If there is a tie between B and C, we favoured C to minimise noise in the final B sample. In the event of a tie among all three positive classes, B was preferred.